Solve the ODE by Linear ODE method:

x dy/dx + 3y = 6x


1
Expert's answer
2021-11-29T19:03:14-0500

x "\frac{dy}{dx}" + 3y = 6x

Expressing the ODE in the form of linear differential equation, "\frac{dy}{dx} + py= q" , where p, q are constants or function of x only we get

"\frac{dy}{dx} + \frac{3}{x}y= 6"

Integrating factor is "e^{\int {\frac{3}{x}dx}}= e^{3ln{x}}=e^{ln{x³}}= x³"

Multiplying both sides by integrating factor , x³

"x³(\frac{dy}{dx} + \frac{3}{x}y)= 6x³"

=> "x³\frac{dy}{dx} + {3x²}y)= 6x³"

=> "x\u00b3dy+ {3x\u00b2}ydx= 6x\u00b3dx"

=> "d{(x\u00b3y)} = 6x\u00b3dx"

Integrating both sides

"\int d{(x³y)} = 6\int x³dx+C'" , where "C'" is integration constant.

=> "x³y = 6. \frac{x⁴}{4} + C'"

=> "x³ y = \frac{3x⁴}{2}+ C'"

=> "2x\u00b3y-3x\u2074 = 2C'"

=> "2x\u00b3y-3x\u2074 = C" where "C = 2C'"

This is the solution of the given differential equation.